# Speak It Out: Solving Symbol-Related Problems with Symbol-to-Language Conversion for Language Models

### 22 Jan 2024 v1

@@ 1 Introduction
Symbols, or more broadly, non-natural language representations such as brackets, digits, molecular formulas, emojis, table delimiters, and abbreviations are ubiquitously encountered in the real world. They are of significant relevance in our daily lives, convey distinct meanings, and play important roles in a variety of tasks such as abstract reasoning (Moskvichev et al., 2023; Xu et al., 2023c), chemical property prediction (Ross et al., 2022; Guo et al., 2023), and tabular question-answering (Chen et al., 2020; Chen, 2023), as exemplified in Table 1. Furthermore, the understanding and reasoning abilities of symbols are of great value in assessing current artificial intelligence (Chollet, 2019).
Large language models (LLMs; Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2022; 2023b; Jiang et al., 2023; Google et al., 2023) have demonstrated impressive abilities in comprehending and generating natural language. GPT-3 (Brown et al., 2020) has showcased capabilities of zero-shot inference that solve problems directly without demonstrations. Kojima et al. (2022) further propose zero-shot-CoT through additional prompting like “Let’s think step by step” to enhance the zero-shot reasoning capability. However, the understanding and reasoning capability of symbols for LLMs still fall behind compared to general natural language. For example, Mitchell et al. (2023) reveal that GPT-4 (OpenAI, 2023b) and GPT-4V (OpenAI, 2023a) only achieve an accuracy of 65% and 25% on minimal abstract reasoning tasks (Moskvichev et al., 2023) requiring inductive reasoning through a series of regular numbers or pixels, which is significantly lower than the human accuracy of 95%. Gendron et al. (2023) further demonstrate that existing LLMs exhibit limited performance for symbol-related problems in contrast with other natural language tasks.
@@
Considering the gap between symbol-based representations and LLMs, our intuition is converting symbols into corresponding language-based expressions. This conversion serves as a bridge to provide LLMs with more friendly and comprehensible information when dealing with symbols. In this paper, we propose symbol-to-language (S2L), a tuning-free method that leverages LLMs to better solve symbol-related problems. The S2L method is both simple and feasible, with its core focus on uncovering language-based expressions that are equivalent or approximate to symbols. In particular, S2L first converts the symbols involved in the problems into their language-based representations, which can either be implemented by prompting the LLMs themselves or by leveraging external tools, such as rules, translators, dictionaries, etc. Then these language-based representations are integrated into the original questions through direct substitution or concatenation, providing valuable contextual information to assist LLMs in solving symbol-related problems.
@@ 4.2 Dyck Language
Remark. The responses during S2L conversion may not be the same among different LLMs. An interesting phenomenon is that GPT-4 recognizes the symbols “<” and “>” as “less than” and “great than”, instead of “open angle bracket” and “close angle bracket”. However, we do not consider this to be a mistake and do not intend to modify the prompts to “correct” the results (e.g., provide some cues indicating these are different types of parentheses). In contrast, we let the LLMs convert the symbols according to their understanding and deduce the final answer through the generated representations by themselves.

### 12 Mar 2024 v2

@@ 1 Introduction
Symbols, or more broadly, non-natural language representations such as brackets, digits, molecular formulas, emojis, table delimiters, and abbreviations are ubiquitously encountered in the real world. They are of significant relevance in our daily lives, convey distinct meanings, and play important roles in a variety of tasks. These symbol-related tasks include abstract reasoning (Moskvichev et al., 2023; Xu et al., 2023c), chemical property prediction (Ross et al., 2022; Guo et al., 2023), and tabular question-answering (Chen et al., 2020; Chen, 2023), as exemplified in Table 1. Consequently, the understanding and reasoning abilities of symbols are of paramount importance for artificial intelligence (Chollet, 2019).
Nowadays, large language models (LLMs; Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2022; 2023b; Jiang et al., 2023; Google et al., 2023) have demonstrated impressive abilities in comprehending and generating natural language. GPT-3 (Brown et al., 2020) has showcased capabilities of zero-shot inference that solve problems directly without demonstrations. Kojima et al. (2022) further propose zero-shot-CoT through additional prompting like “Let’s think step by step” to enhance the zero-shot reasoning capability. However, the understanding and reasoning capability of symbols for LLMs still fall behind compared to general natural language. For example, Mitchell et al. (2023) reveal that GPT-4 (OpenAI, 2023b) and GPT-4V (OpenAI, 2023a) only achieve an accuracy of 65% and 25% on minimal abstract reasoning tasks (Moskvichev et al., 2023) requiring inductive reasoning through a series of regular numbers or pixels, which is significantly lower than the human accuracy of 95%. Gendron et al. (2023) further demonstrate that existing LLMs exhibit limited performance for symbol-related problems in contrast with other natural language tasks.
@@
Considering the gap between symbol-based representations and LLMs, our intuition is converting symbols into corresponding language-based expressions. This conversion serves as a bridge, providing LLMs with more friendly and comprehensible information. In this paper, we propose symbol-to-language (S2L), a tuning-free method that leverages LLMs to better solve symbol-related problems. The S2L method is both simple and feasible, with its core focus on uncovering language-based expressions that are equivalent or approximate to symbols. In particular, S2L first converts the symbols involved in the problems into their language-based representations, which can either be implemented by prompting the LLMs themselves or by leveraging external cost-friendly tools, such as rules, translators, dictionaries, etc. Then these language-based representations are integrated into the original questions through direct substitution or concatenation, providing valuable contextual information to assist LLMs in solving symbol-related problems.
@@ 4 Experiments
(Zero-Shot) | |||||||
gpt-4 | |||||||
+S2L w/ model | () | () | () | () | () | () | () |
+S2L w/ rule | () | () | () | () | () | () | () |
gpt-3.5-turbo | |||||||
+S2L w/ model | () | () | () | () | () | () | () |
+S2L w/ rule | () | () | () | () | () | () | () |
openchat-3.5-7b | |||||||
+S2L w/ model | () | () | () | () | () | () | () |
+S2L w/ rule | () | () | () | () | () | () | () |
(Zero-Shot-CoT) | |||||||
@@ 4.1 Abstract Reasoning
(Zero-Shot-CoT) | ||||||
gpt-4 | ||||||
+S2L w/ model | () | () | () | () | () | () |
gpt-3.5-turbo | ||||||
+S2L w/ model | () | () | () | () | () | () |
openchat-3.5-7b | ||||||
+S2L w/ model | () | () | () | () | () | () |
@@ 4.2 Dyck Language
Remark. The responses during S2L conversion may not be the same among different LLMs. An interesting phenomenon is that GPT-4 recognizes the symbols “<” and “>” as “less than” and “greater than”, instead of “open angle bracket” and “close angle bracket”. However, we do not consider this to be a mistake and do not intend to modify the prompts to “correct” the results (e.g., provide some cues indicating these are different types of parentheses). In contrast, we let the LLMs convert the symbols according to their understanding and deduce the final answer through the generated representations by themselves.
@@ 6 Conclusion
Appendix A Case Study
We show some examples in Figures 8, 9, 10, 11, 12, and 13 where our symbol-to-language gives correct responses using the GPT-4 model. Please refer to https://github.com/THUNLP-MT/symbol2language for detailed results.